Grade 12th passAlgebra

Convert each part of the ratio to percentage:
1:4 ////

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to find total parts by adding the given ratio = 1 + 4 = 5
1stpart = (1/5) = (1/5) × 100 %
= 1 × 20%
= 20%
2ndpart = (4/5) = (4/5) × 100%
= 4 × 20 = 80%

Thanks